Please help me compute the last parts of required 3,4,5.

Factor Company is planning to add a new product to its line. To manufacture this product, the company needs to buy a new machine at a $480,000 cost with an expected four-year life and a $20,000 salvage value. All sales are for cash, and all costs are out-of-pocket, except for depreciation on the new machine. Additional information includes the following. (PV of $1, FV of $1, PVA of $1, and FVA of $1) (Use appropriate factor(s) from the tables provided.) Expected annual sales of new product Expected annual costs of new product $1,840,000 Direct materials Direct labor overhead (excluding straight-line depreciation on new machine) Selling and administrative expenses Income taxes 480,000 672,000 336,000 160,000 30% Required: 1. Compute straight-line depreciation for each year of this new machines life. 2. Determine expected net income and net cash flow for each year of this machines life. 3. Compute this machines payback period, assuming that cash flows occur evenly throughout each year. 4. Compute this machines accounting rate of return, assuming that income is earned evenly throughout each year. 5. Compute the net present value for this machine using a discount rate of 7% and assuming that cash flows occur at each year-end. Hint: Salvage value is a cash inflow at the end of the assets life.)
Answer #1

3)Payback period = cost of investment / annual net cash flow

       = 480000/168900

        = 2.84 years

4)accounting rate of return = Net Income / annual average investment

         = 53900 / 250000

          = .2156 or 21.56%

Average investment = [beginning book value +ending book value ]/2

    =[480000+20000]/2

= 250000

5)Present value of cash inflow = [PVA7%,4 *Annual cash inflow ]+[PVF 7%,4*Salvage value]

        =[3.38721*168900]+ [.76290*20000]

       = 572099.77+ 15258

         = 587357.77

NPV =Present value of cash inflow -present value of cash outflow

                   = 587357.77- 480000

                    = 107357.77